OpenStudy (anonymous):

rewrite the following as a base-ten number 6*10^6+2*10^4+5*10^3

OpenStudy (anonymous):

and what wold be the base-ten number? 620,500

OpenStudy (unklerhaukus):

yeah i think i answered the wrong question by mistake

OpenStudy (unklerhaukus):

\[6\times10^6+2\times10^4+5\times10^3=6000000+20000+5000=602500=6.025\times 10^6\]

OpenStudy (unklerhaukus):

\[=6.025\times 10^6\]

OpenStudy (anonymous):

thanks anyway but the answer was 6,025,000